OpenStudy (anonymous):

If (a+b)^2=9 and (a-b)^2=49 what is the value of a^2+b^2 ?

OpenStudy (dape):

\[ \frac{(a+b)^2+(a-b)^2}{2} = a^2+b^2 \] \[ a^2+b^2=\frac{9+49}{2}=29 \]

OpenStudy (anonymous):

Thank you